OpenStudy (anonymous):

A 20-cm long and 1.0-mm diameter thick copper wire (resistivity 1.7 × 10-8 Ωm) is connected across a 3-V battery. The current through the wire is

OpenStudy (anonymous):

Could someone explain. I am having a very hard time with physics. I know this may be simple to a lot of people but I'm lost.

OpenStudy (ash2326):

We know from Ohm's law, that current is \[I=\frac{V}{R}\] we have v voltage given as 3 volts we need to find R resistance We know \[R=\rho \frac l A\] " 20-cm long and 1.0-mm diameter thick copper wire (resistivity 1.7 × 10-8 Ωm)" length= 20 cm or 20 x10^(-2) meters Area= \[\pi r^2\] We have diameter as 1 mm so radius 0.5 mm \[A=\pi r^2=\pi (0.5\times 10^{-3})^2\] Resistivity \(\rho\) is given \[\rho=1.7 \times 10^{-8}\]

OpenStudy (ash2326):

First find resistance , then you can find current. Let me know if you have any trouble